OpenStudy (anonymous):

Nathan and Steven stacked boxes on a shelf. Nathan lifted 13 boxes and Steven lifted 15 boxes. The boxes that Nathan lifted each weighed 12 lb more than the boxes Steven lifted. Which expression represents the total number of pounds Nathan lifted? Let s represent the weight of the boxes that Steven lifted. A. 13(s +12) B. s + 12 C. 15(s + 12) D. 12(s + 13)

OpenStudy (anonymous):

help

OpenStudy (anonymous):

It would be A.

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Two stores sell the same computer for the same original price. Store A advertises that the computer is on sale for 25% off the original price. Store B advertises that it is reducing the computer’s price by $180. When Brittany compares the sale prices of the computer in both stores, she concludes that the sale prices are equal. Which equation models this situation? Let p represent the computer’s original price. A. 0.25p = p – 180 B. 0.75p = p + 180 C. 0.25(p – 180) = p D. 0.75p = p − 180